AB INITIO POTENTIAL ENERGY SURFACES FOR H_{5}^{+} AND CH_{5}^{+}

\underline{JOEL M. BOWMAN}, ZHEN XIE, ZHONG JIN and BASTIAAN BRAAMS, Department of Chemistry and Cherry L. Emerson Center for Scientific Computation, Emory University, Atlanta, GA 30322.

I will describe new ab initio potential energy surfaces for H_{5}^{+} and CH_{5}^{+} and present some results of vibrational and scattering calculations using these surfaces.